Session: Feb 1999 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K.Borkowski at Thu Feb 18 09:43:52 MET 1999) Success
No problems to report.

Onsala (A. Polatidis at Thu Feb 18 10:52:11 MET 1999) Success

Medicina (G.Maccaferri at Thu Feb 18 15:05:14 MET 1999) Success

Jodrell (Alastair Gunn at Fri Feb 19 14:17:50 MET 1999) Minor failures
Receiver began to warm sometime before 23:00 UT. Attenuation levels were reset but data probably bad for rest of experiment. Receiver was replaced after this experiment.

Cambridge (Alastair Gunn at Fri Feb 19 14:18:22 MET 1999) Minor failures
Upconverion LO value had to be reset within the first hour, was okay at start but was mysteriously changed. Possibly some data loss within first hour. No other known problems.

Shanghai (Huang Xinyong at Mon Feb 22 10:58:16 MET 1999) Success
It is raining.

Effelsberg (dgraham at Mon Feb 22 15:33:21 MET 1999) Failure
stop for snow 1851-1919,2030-2159,0039-0551, so most of experiment lost.

Westerbork (Tony Foley at Mon Feb 22 21:16:57 MET 1999) Success
No known problems

Noto (C.Trigilio at Mon Feb 22 23:24:21 MET 1999) Success

Urumqi (Liu Xiang at Tue Feb 23 05:12:27 MET 1999) Success

Calibration (Jean\-Francois Desmurs at Wed Mar 3 11:22:37 MET 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the feb99 directory under the name: em021a.antab. Read the comments in the file to get further information. (note: Ef Tsys wrong during part of the time; Ur missed).